OpenStudy (anonymous):

3 senior tickets and 1 child ticket =38$ next day.. 3 senior tickets & 2 child tickets = 52$ find the price of a senior citizen ticket and the price of the child ticket

OpenStudy (anonymous):

senior tickets=8$ and child tickets=14$

OpenStudy (anonymous):

let s be the price of a senior ticket, and c be the price of a child ticket. 3s + c = 38 3s + 2c = 52. Subtracting the first equation from the second we get c = 14 We then substitute the value for c into one of the equations to solve for s. 3s +(14) = 38 3s = 24 s = 8. A child's ticket is $14, and a senior's ticket is $8

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

@ manucub0 how did you get 14?

OpenStudy (anonymous):

subtract equation 2 from 1 3s +c=38 -(3s+2c=52) this will give -c=-14 which is c=14

OpenStudy (anonymous):

i mean how did you simplifyt that :(

OpenStudy (anonymous):

well change the sign of evry term in second equation and then just add like u add two simple equations. and u will get that

OpenStudy (anonymous):

can you show me step by step :(

OpenStudy (anonymous):

ok : from ur problem : on first day there were 3 seniors and 1 child . lets denote seniors by s and children by c . cost of first day was $38 so it mens that : 3s+c=38 ----eq(1) similarly for second day : 3s+2c=52 ---eq(2) now to solve equations in 2-variables system we try to remove one variable . so as can be observed 3s is there in both equations so if we subtract eq(2) from eq(1) then terms containing s will cancel out . u can also do eq(2) -eq(1) so eq(2)-eq(1) gives : -c=-14 this gives c=14 now we put this value in any of the equations so putting c=14 in eq(1) 3s+c=38 or 3s+14=38 3s=38-14=24 3s=24 gives s=8 so we have now senior cost =$8 denoted by s child cost =$14 denoted by c i hope u are clear now ..